OpenStudy (anonymous):

What are the zeroes of f(x)= x^2 +5x+6? A- x= -2, -3 B- x=2,3 C- x= -2,3 D- = 2,-3

OpenStudy (anonymous):

Well we need to find the factors of the polynomial

OpenStudy (michele_laino):

hint: we have to solve the subsequent quadratic equation: \[\Large {x^2} + 5x + 6 = 0\]

OpenStudy (michele_laino):

please apply the standard formula: \[\Large \begin{gathered} x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}} \hfill \\ \\ a = 1,\;b = 5,\;c = 6 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

we can either use use the quadratic equation, complete the square or any other ways of solving polynomials.

OpenStudy (anonymous):

I'm honestly so lost @deeezzzzz